OpenStudy (anonymous):

Edison has a dog walking business. He earns $6.00 for each dog that he walks. Edison walked 4 dogs on Friday, 6 dogs on Saturday, and some additional dogs on Sunday. He earned a total of $108.00 for those three days. How many dogs did he walk on Sunday?

OpenStudy (anonymous):

the answer is 8 NVM

OpenStudy (anonymous):

Let's call the total number of dogs 'D". Let's call the number of dogs on Friday 'f', the number of dogs on Saturday 's' and the number of dogs on Sunday 'u'. We are told that \[108=6f+6s+6u\] or in simplest terms: \[108=6(f+s+u)\] WE are also told that \[f=4\] and\[s=6\] So: \[108=6(4+6+u)\]\[108=6(u+10)\]\[18=u+10\]\[u=8\] You can check by substitution: \[108=6(4+6+8)\]\[108=108\] Checks.